OpenStudy (anonymous):

Solve. (x - 5)^2 = 4.

OpenStudy (anonymous):

(x-5)(x-5)=4 x^2-10x+25=4 x^2-10x+21=0 (x-7)(x-3)=0 x=7, x=3